Multiply:
(5√[2] + 3)(√[2] - 2√[3])
Please show howyou get the answer??

Answer by sdmmadam@yahoo.com(530)

Multiply:
(5√[2] + 3)(√[2] - 2√[3])
=5√[2](√[2] - 2√[3]) + 3X(√[2] - 2√[3])
[using (a+b)(c+d) = a(c+d) + b(c+d) = (ac+ad)+(bc+bd), first term(along with the sign) of the first bracket multiplied with the whole of the next bracket plus the second term (along with the sign)of the first bracket multiplied with the whole of the next bracket]
=5√[2]X(√[2] - 5√[2]X 2√[3] + 3√[2] - 6√[3])
=5X(2)-10√[6]+ 3√[2] - 6√[3]) (we have √[a](√[b] = √[(ab)] )
=10-10√[6]+ 3√[2] - 6√[3])